What is wrong with the solution of the student?
x(x+3) = 9
x = 9 or x+3 = 9
x = 6
So the roots are 9 and 6.

Answer by mananth(16946)

x(x+3)=9
x^2+3x=9=0
x^2+3x-9=0
a=1,b=3,c=-9
x1= -3 +sqrt(9+36)/2
x1=-3+sqrt45/2
x1= 1.85
...
x2= -4.85